# Non-Linear Model-Based Sequential Decision-Making in Agriculture

### v1

@@ 3.3 Nonlinear model-based UCB
The UCB (Upper Confidence Bound) algorithm (Auer and Ortner,, 2010; Chu et al.,, 2011; Zhou et al.,, 2020) embodies the principle of “optimism in the face of uncertainty” by selecting actions that maximize an upper confidence bound on the expected reward. In each round, it favors actions that either appear promising based on current estimates or for which the uncertainty is high, thereby systematically encouraging exploration as well as exploitation.
@@
UCB algorithms provide strong theoretical guarantees for balancing exploration and exploitation, and are particularly effective when reliable, model-based estimates of prediction uncertainty are available (see, e.g., (Zhou et al.,, 2020)). It is particularly advantageous in settings where data collection is expensive or risky, as it prioritizes actions that could yield high rewards or that have not been thoroughly explored.
Algorithm 3 UCB for Nonlinear Contextual Bandits
@@ 3.4 ViOlin (Virtual Ascent with Online Model Learner)
The ViOlin algorithm (Algorithm 4(Dong et al.,, 2021) is a model-based bandit strategy designed for efficient learning in nonlinear settings. Unlike $\epsilon$-greedy, which explores actions uniformly at random, or UCB, which prioritizes actions with high uncertainty, ViOlin is a greedy method that leverages both the current model estimate and local geometric information about the reward surface (i.e., gradient and curvature). At each round, the algorithm chooses the action predicted to be the best for profit maximization by the current estimated model, then it updates the model using not only the observed rewards, but also estimates of the local gradient and Hessian of the reward function at the selected action. This curvature-matching approach yields strong sample efficiency guarantees, especially in complex, high-dimensional, or highly nonlinear problems, by ensuring the model learns both the value and the local shape of the reward function. While originally, Dong et al., (2021) present the ViOlin algorithm for the non-noisy setting, we illustrate its performance under model noise as well.
@@
ViOlin is particularly attractive in data-limited settings (e.g., smallholder or resource-constrained agricultural experiments), where minimizing the number of field trials is crucial. ViOlin’s greedy exploration is guided by the model fit and its local geometry, without explicit randomized or uncertainty-driven action selection. This enables fast convergence to (local) optima when the model class matches the true reward structure and curvature information is reliable. For well-behaved, low-dimensional agronomic models, the performance of ViOlin may be comparable to that of simpler methods such as $\epsilon$-greedy and UCB. However, as model complexity such as multi-modal response curves or as dimensionality of $\mathcal{X}$ increases, ViOlin’s sample efficiency guarantees make it especially advantageous. This makes ViOlin well suited for adaptive experimentation in limited data regimes, such as on-farm field trials or precision agriculture for smallholder contexts, where each data point is costly to obtain.
@@
While the primary emphasis of this work is not on theoretical development, we present and discuss several well-known theoretical results and highlight their implications for the nonlinear model-based contextual bandit framework.
@@ 4 Theoretical guarantees for nonlinear bandit algorithms
In this section, we characterize the regret performance of sequential algorithms under increasingly flexible model-based reward mechanisms, beginning with classical linear models, then considering nonparametric function classes, nonlinear parametric (mechanistic) models as used in agronomy, and finally accounting for the impact of model misspecification.
Minimax regret for linear and GLM bandits.
@@
where $T$ is the time horizon (Abbasi-Yadkori et al.,, 2011; Dani et al.,, 2008). Generalized linear bandit models achieve similar rates up to logarithmic factors (Filippi et al.,, 2010; Bastani et al.,, 2020).
@@
For more general nonparametric function classes, such as Lipschitz or RKHS (kernelized) reward functions, the minimax regret scales polynomially with the dimension $d$ of the action/context space (Slivkins,, 2014; Srinivas et al.,, 2012):
$R_{T}=\widetilde{O}\left(T^{\frac{d+1}{d+2}}\right),$
where $d$ is the relevant dimension. These models are highly flexible but suffer from poor sample efficiency in high dimensions, motivating the use of structured parametric models when such domain knowledge is available.
NeuralUCB and nonlinear function classes.
Recent advances have extended bandit algorithms to rich nonlinear function classes via neural networks. In particular, the NeuralUCB algorithm (Zhou et al.,, 2020) achieves a regret bound of
@@
where $\tilde{d}(\lambda)$ denotes the effective dimension of the neural tangent kernel (NTK) associated with the underlying neural network, and $\widetilde{O}$ suppresses logarithmic dependencies on $T$ and $1/\lambda$. The effective dimension is defined as
$\tilde{d}(\lambda)=\mathrm{Tr}\left(K_{\mathcal{X},\mathcal{X}}(K_{\mathcal{X},\mathcal{X}}+\lambda I)^{-1}\right),$
where $K_{\mathcal{X},\mathcal{X}}$ is the NTK Gram matrix over the sequence of played arms $\{x_{1},\dots,x_{T}\}$ and $\lambda>0$ is a regularization parameter. For a two-layer ReLU neural network with input dimension $d$ and width $m$, it holds that $\tilde{d}(\lambda)\leq d\,\mathrm{poly}(m,\log(T/\lambda))$, so that the regret rate can be written as
@@
Here, $\mathrm{poly}(m,\cdot)$ denotes a polynomial function of the network width and the indicated arguments. Thus, the regret bound increases with both the expressiveness of the neural network (through $m$) and the ambient input dimension $d$, but retains the $\sqrt{T}$ scaling characteristic of parametric bandit models. This result demonstrates that, under suitable overparameterization, nonlinear bandits with neural function approximation can achieve regret scaling close to the linear case (up to polynomial factors in $m$), substantially improving sample efficiency over generic nonparametric rates. However, this improvement is obtained at the cost of increased computational complexity and model tuning.
Table 1: Summary of Regret Rates for Contextual Bandits under Different Model Classes
@@
Building on these insights, we now focus on model-based nonlinear bandits, a regime particularly relevant in agronomy, where biological processes often admit interpretable, low-dimensional parameterizations. Since one of our motivations for this work is to provide decision-making algorithms in sample-limited scenarios, we focus on characterizing sample complexity more than providing regret guarantees in this section. Recent work by Dong et al., (2021) has shown a breakthrough for model-based nonlinear bandits: for any class of reward functions with bounded sequential Rademacher complexity, it is possible to find an $\epsilon$-approximate local maximum with sample complexity polynomial in the complexity of the model class, independent of the action dimension. This suggests that, unlike classical methods whose sample complexity or regret is exponential in dimension, model-based nonlinear bandits can be vastly more efficient when the reward function admits a suitable low-complexity parameterization.
Sample Complexity and Local Regret for Nonlinear Bandits.
In the nonlinear bandit setting, the goal is not always to find a global maximum (which is often computationally and statistically intractable), but rather to identify an approximate local maximum of the unknown reward function as efficiently as possible. Here, sample complexity refers to the minimum number of experimental rounds required to guarantee, with high probability, that the recommended action achieves performance within a specified tolerance of a local optimum. This guarantee is formalized through the notion of local regret, which measures the suboptimality of the chosen action relative to the best locally optimal action. In practice, for well-behaved (e.g., unimodal) functions such as those common in agronomic yield response, a local maximum often coincides with the global maximum.
@@
We defer the proof of Theorem 2 to the Appendix (Section A) for brevity and discuss its implications from an application viewpoint. Our theoretical result complement Dong et al., (2021)’s general framework by concretely instantiating the sample complexity for these practically important nonlinear classes, including quadratic plateau, Michaelis-Menten, Mitscherlich, and logistic models—the sequential Rademacher complexity $\mathfrak{R}_{T}^{\mathrm{seq}}(\mathcal{F})$ can be bounded as $\widetilde{O}(1/\sqrt{T})$. Therefore, the sample complexity for identifying an $\epsilon$-optimal arm (in the sense of a local maximum) in these bandit settings scales as $\widetilde{O}(1/\epsilon^{8})$, with constants depending on the parameter and input bounds. Note that, this result of Dong et al., (2021) is not tight and can potentially be improved in terms of the polynomial dependence on $\epsilon$.
@@
While Dong et al., (2021) establish general sample complexity and regret guarantees for nonlinear bandit problems for their proposed algorithm (Algorithm 4) including highly complex model classes like 2-layer Neural networks, our agricultural decision models are low-dimensional and interpretable (e.g., quadratic threshold). In this favorable case, classic algorithms such as UCB and $\epsilon$-greedy are not only interpretable and easy to implement but also benefit from even stronger theoretical guarantees than those stated in Dong et al., (2021), due to the reduced complexity of the model class. Therefore, we use these classic model-based approaches and provide sample complexity statements derived from Dong et al., (2021) as a conservative bound.
@@
Most contextual bandit algorithms rely on the assumption of a well-specified model (or “realizability”), that is, the existence of a function $f^{*}\in\mathcal{F}$ such that the mean reward is $\mu(x)=f^{*}(x)$ for all actions $x$ (Chu et al.,, 2011; Abbasi-Yadkori et al.,, 2011; Agarwal et al.,, 2012; Foster et al.,, 2018). However, in practice, exact realizability rarely holds, and it is important to consider the effect of model misspecification.
Recent work (see, e.g., (Greenewald et al.,, 2021; Lattimore et al.,, 2020; Ghosh et al.,, 2017)) has studied contextual bandits under various notions of misspecification. A common formulation is the uniform $\varepsilon$-misspecification setting, where
@@
For linear contextual bandits, Lattimore et al., (2020) show that the cumulative regret must satisfy
@@
5 Simulation study: Emulating real-world agricultural decision-making
@@ 5.1 From yield to profit: Economic optimization of fertilizer
    Model-based $\epsilon$-greedy algorithm (Algorithm 5): This method selects a random fertilizer rate with probability $\epsilon_{t}=t^{-a}$, with $a$ chosen from $\{0.5,1,1.5\}$, for exploration. Otherwise, it chooses the rate closest to the current profit-maximizing estimate $x^{*}$, computed using the fitted model and closed-form expressions where available.
@@
Although these methods differ in their exploration strategies, they all share the same structure: fitting a nonlinear model to guide adaptive decision-making under uncertainty. Accurate initialization of nonlinear model parameters is crucial to the success of these algorithms, especially in small-sample settings. Poor initial values can mislead early optimization steps and delay convergence. To benchmark the performance of nonlinear model-based strategies, we compare them against two widely used baselines: LinUCB (Chu et al.,, 2011) and $k$-Nearest Neighbor-UCB (kNN-UCB) (Reeve et al.,, 2018).
@@
Table 2: Closed-form expressions for the profit-maximizing fertilizer dose $x^{*}$ for different yield response models.
@@ 5.2 Well-specified setting
In the well-specified setting, data were generated according to a quadratic threshold model of the form:
@@
We evaluated five algorithms: model-based $\epsilon$-greedy, nonlinear UCB, ViOlin (curvature-matched nonlinear bandit), LinUCB (linear model-based UCB), and kNN-UCB (nonparametric UCB using $k=3$ nearest neighbors). Hyperparameters for each of the algorithms are chosen based on theoretical guidelines and grid searches over multiple runs. We choose $\epsilon_{t}=t^{-0.5}$ for $\epsilon$-greedy in this setting, $\alpha=1$ in UCB, linUCB and kNN-UCB for uniformity and fairness in comparison, $\alpha_{1}=2.0,\alpha_{2}=640$ for the ViOlin algorithm based on guidelines in Dong et al., (2021). For each algorithm and parameter configuration, we performed $10$ independent simulation replicates of $T=30$ rounds. Each algorithm fits its specified model to the observed data and selects actions to maximize estimated profit. The regret at each round is defined as the difference between the average profit for the optimal arm (computed using knowledge of the true reward function) and the average profit realized by the arm selected by the algorithm, that is,
@@
where $x^{*}$ denotes the arm maximizing the true profit function, $\theta^{*}$ the true model parameters and $x_{t}$ is the arm selected at round $t$ by the bandit algorithm employed to make decisions. For each run, we recorded cumulative regret, average profit, and selected fertilizer rates.
@@
Figure 1 presents the average cumulative regret and the average profit per round for each algorithm. Several trends are immediately apparent:
@@
Similar trends were observed across other values of the fertilizer price $p_{x}$ and for all nonlinear response models considered. Taken together, these simulation studies demonstrate a consistent advantage of model-based nonlinear bandit algorithms over both nonparametric and linear parametric alternatives in small-sample regimes. Based on these findings, we subsequently focus our comparisons on the family of nonlinear model-based algorithms, investigating their relative strengths under a range of simulation scenarios to clarify which approaches are most advantageous in which settings.
Figure 1: Comparing different Bandit algorithms for profit maximization/regret minimization
@@
In this next part of the simulation study, we focus on illustrating the implications of using the three non-linear model-based bandit algorithms for learning fertilizer rates sequentially with the goal of profit maximization. Again in a small-sample regime, we run each of our algorithms for $T=30$ rounds and replicate each run 10 times. Since our algorithms choose arms over time, it is important to visualize how the arm choices evolve over time for each of the algorithms. In Figure 2(a), we plot the running proportion of fertilizer rate ranging from $\{0,50,\ldots,250\}$ lb N/ac, selected at each round by each of the algorithm over the 10 replicates for the quadratic plateau model for $p_{x}=\$0.5$/lb N to mimic realistic prices per unit nitrogen fertilizer (Urea) in the Midwest US. Note that in the beginning, the UCB and $\epsilon$-greedy ($\epsilon_{t}=t^{-1.5}$) explore other arms such as $x=100,150$ lb N/ac, but just after a few rounds of exploration, learn the better arm to be $x=200$ lb N/ac. ViOlin is more greedy from the beginning itself as can be seen by the proportions of other arms (other than purple) selected in the initial rounds. In Figure 2(b), in order to assess the affect of increasing fertilizer prices on decision-making for profit maximization, we also plotted the average selected fertilizer rate over time for the three algorithms. We note that the optimum arm (fertilizer rate) choice decreases as $p_{x}$ (price per unit) increases from \$0.3 to \$0.7/lb N. Additionally, since our decisions depend significantly on how well our parameters for the non-linear model are estimated over time, in Figure 3, we also plot the parameter trajectories over time for the Quadratic Threshold model with $p_{X}=\$0.5$/lb N. Note that all the three algorithms, as data accumulates, the parameter estimates stabilize close to the true parameter values that were used to generate the data. In terms of interpretability, one can, for example, look at the parameter estimates for $x_{0}$ (bottom right), which denotes the threshold value at which the quadratic model plateaus, thus reflecting the saturating point over rounds (or seasons), beyond which additional fertilizer amount provide little benefit. We also checked for sensitivity to initial values and we noticed that good initial values matter for better convergence to the true parameters, especially in a small sample regime. Therefore, we recommend using domain-knowledge to set initial values of the parameters in the least-squares estimation algorithm. We conducted similar experiments for the other non-linear models and those results are presented in the Appendix.
Figure 3: Parameter trajectories for the quadratic threshold model over rounds

### v2

@@ 3.3 Nonlinear model-based UCB
The UCB (Upper Confidence Bound) algorithm (Auer and Ortner,, 2010; Chu et al.,, 2011; Zhou et al.,, 2020) embodies the principle of “optimism in the face of uncertainty” by selecting actions that maximize an upper confidence bound on the expected reward. Here, the nonlinear model is repeatedly fitted, and for each input level the algorithm constructs an upper confidence bound on the predicted reward. The next action is chosen to maximize this upper bound. This mechanism naturally prioritizes inputs that are either promising in terms of mean yield or remain highly uncertain, thereby encouraging exploration in scientifically justified directions. Applied to fertilizer management, UCB corresponds to testing input levels where either predicted yield is high or parameter uncertainty remains large.
@@
UCB algorithms provide strong theoretical guarantees for balancing exploration and exploitation, and are particularly effective when reliable, model-based estimates of prediction uncertainty are available (see, e.g., (Zhou et al.,, 2020)). It is particularly advantageous in settings where data collection is expensive or risky, as it prioritizes actions that could yield high rewards or that have not been thoroughly explored. Model-based UCB (Algorithm 6) thus combines statistical rigor with the agronomic interpretability of nonlinear functions, making it a strong candidate for adaptive on-farm experimentation.
Algorithm 3 Nonlinear model-based UCB
@@ 3.4 ViOlin (Virtual Ascent with Online Model Learner)
The ViOlin algorithm (Algorithm 4(Dong et al.,, 2021) is a model-based bandit strategy designed for efficient learning in nonlinear settings. Unlike $\epsilon$-greedy, which explores actions uniformly at random, or UCB, which prioritizes actions with high uncertainty, ViOlin is a greedy method that leverages both the current model estimate and local geometric information about the reward surface (i.e., gradient and curvature). At each round, the algorithm chooses the action predicted to be the best for profit maximization by the current estimated model, then it updates the model using not only the observed rewards, but also estimates of the local gradient and curvature of the reward function at the selected action. This curvature-matching approach yields strong sample efficiency guarantees, especially in complex, high-dimensional, or highly nonlinear problems, by ensuring the model learns both the value and the local shape of the reward function. While originally, Dong et al., (2021) present the ViOlin algorithm for the non-noisy setting, we illustrate its performance under model noise as well.
@@
ViOlin is particularly attractive in data-limited settings (e.g., smallholder or resource-constrained agricultural experiments), where minimizing the number of field trials is crucial. ViOlin’s greedy exploration is guided by the model fit and its local geometry, without explicit randomized or uncertainty-driven action selection. This enables fast convergence to (local) optima when the model class matches the true reward structure and curvature information is reliable. In low-dimensional agronomic models with smooth response surfaces, ViOlin performs on par with simpler strategies such as $\epsilon$-greedy and UCB. Its strength becomes pronounced as complexity grows: in multi-modal settings or higher-dimensional decision spaces, ViOlin’s sample-efficiency guarantees provide a clear advantage. This makes ViOlin well suited for adaptive experimentation in limited data regimes, such as on-farm field trials or precision agriculture for smallholder contexts, where each data point is costly to obtain.
@@
To illustrate how these algorithms operate and compare in practice, we include a simple, step-by-step example using the Mitscherlich model over three or four rounds in Section 3 of the Appendix. This concrete illustration should help readers, especially those less familiar with bandit algorithms, understand the exploration-exploitation tradeoff.
While the primary emphasis of this work is not on theoretical development, we present and discuss several well-known theoretical results and highlight their implications for the nonlinear model-based bandit framework.
@@ 4 Theoretical guarantees for nonlinear bandit algorithms
In this section, we characterize the regret and sample-complexity guarantees of sequential algorithms under increasingly flexible reward models, beginning with classical linear formulations, extending to nonparametric function classes, then focusing on nonlinear parametric (mechanistic) models as used in agronomy, and finally accounting for model misspecification. While the general theory applies broadly, our emphasis is on sample efficiency and robustness in low-data agricultural regimes, where each field trial is costly and decision errors translate to both economic and environmental consequences.
Linear and GLM bandits.
@@
where $T$ is the time horizon (Abbasi-Yadkori et al.,, 2011; Dani et al.,, 2008). Generalized linear bandit models achieve similar rates up to logarithmic factors (Filippi et al.,, 2010; Bastani et al.,, 2020). These results provide a baseline: when yield or profit responds linearly to inputs, bandit algorithms can achieve fast learning rates. However, such linear approximations are rarely biologically realistic in crop response.
@@
For more general nonparametric function classes, such as Lipschitz or RKHS (kernelized) reward functions, the minimax regret scales polynomially with the dimension $d$ of the action/context space $\mathcal{X}$  (Slivkins,, 2014; Srinivas et al.,, 2012):
$R_{T}=\widetilde{O}\left(T^{\frac{d+1}{d+2}}\right).$
While attractive for flexibility, these guarantees deteriorate with dimension and are impractical when only a small number of seasons or trials are available, which is a common reality in agriculture.
NeuralUCB and expressive function classes.
Recent work has extended bandit algorithms to highly expressive nonlinear function classes via neural networks. For instance, NeuralUCB (Zhou et al.,, 2020) achieves regret rates of order
@@
where $\tilde{d}(\lambda)$ is the effective dimension of the neural tangent kernel associated with the underlying network. For example, for a two-layer ReLU neural network with input dimension $d$ and width $m$, it holds that $\tilde{d}(\lambda)\leq d\,\mathrm{poly}(m,\log(T/\lambda))$, so that the regret rate can be written as
@@
Here, $\mathrm{poly}(m,\cdot)$ denotes a polynomial function of the network width and the indicated arguments. Thus, the regret bound increases with both the expressiveness of the neural network (through $m$) and the ambient input dimension $d$, but retains the $\sqrt{T}$ scaling characteristic of parametric (linear) bandit models. However, these improvements come with heavy computational costs, the need for large data, and limited interpretability.
We summarize the regret bounds for the four classes of bandit problems in Table 1. While such results highlight the theoretical reach of modern methods, they are poorly suited to agricultural decision-making, where data are scarce, feedback is seasonal, and recommendations must be transparent to practitioners. Our focus, therefore, is on mechanistic yield-response models such as Mitscherlich or Michaelis-Menten. These models are parsimonious, interpretable, and biologically grounded, yet flexible enough to capture crop dose-response behavior. This specialization ensures theoretical guarantees translate into actionable insights, directly relevant for resource-efficient agricultural experimentation.
Table 1: Summary of Regret Rates for Bandits under Different Model Classes
@@
Building on these insights, we now focus on model-based nonlinear bandits, a regime particularly relevant in agronomy, where biological processes often admit interpretable, low-dimensional parameterizations. Since one of our motivations for this work is to provide decision-making algorithms in sample-limited scenarios, we focus on characterizing sample complexity more than providing regret guarantees in this section. Recent work by Dong et al., (2021) has shown a breakthrough for model-based nonlinear bandits: for any class of reward functions with bounded sequential Rademacher complexity, it is possible to find an $\epsilon$-approximate local maximum with sample complexity polynomial in the complexity of the model class, independent of the action dimension. This suggests that, unlike classical methods whose sample complexity or regret is exponential in dimension, model-based methods can be vastly more efficient when the reward function admits a suitable low-complexity parameterization.
Sample Complexity and Local Regret for Model-based Nonlinear Bandits.
In the model-based bandit setting, the functional form of the reward is known, but its parameters are not. The goal is therefore to efficiently identify an approximate local maximum of this parametric reward function, rather than a global maximum which is often computationally and statistically intractable. Here, sample complexity refers to the minimum number of experimental rounds required to guarantee, with high probability, that the recommended action achieves performance within a specified tolerance of a local optimum. This guarantee is formalized through the notion of local regret, which measures the suboptimality of the chosen action relative to the best locally optimal action. In practice, for well-behaved (e.g., unimodal) functions such as those common in agronomic yield response, a local maximum often coincides with the global maximum.
@@
We defer the proof of Theorem 2 to the Appendix (Section A) for brevity and discuss its implications from an application viewpoint. Our theoretical result complement Dong et al., (2021)’s general framework by concretely instantiating the sample complexity for these practically important nonlinear classes, including quadratic plateau (threshold), Michaelis-Menten, Mitscherlich, and logistic models, the sequential Rademacher complexity $\mathfrak{R}_{T}^{\mathrm{seq}}(\mathcal{F})$ can be bounded as $\widetilde{O}(1/\sqrt{T})$. Therefore, the sample complexity for identifying an $\epsilon$-optimal arm (in the sense of a local maximum) in these bandit settings scales as $\widetilde{O}(1/\epsilon^{8})$, with constants depending on the parameter and input bounds. This bound, as established in Dong et al., (2021), is general and not specialized to the structure of agronomic models in terms of its polynomial dependence on $\epsilon$. Deriving sharper rates for the structured nonlinear models considered here is an important direction for future work.
@@
Dong et al., (2021) provide broad sample complexity and regret guarantees for nonlinear bandit problems, including highly expressive classes such as 2-layer neural networks. By contrast, our agricultural decision models are low-dimensional and mechanistic (e.g., quadratic plateau), which makes them both interpretable and practically relevant. In this structured setting, algorithms such as UCB and $\epsilon$-greedy are not only straightforward to implement but also expected to admit stronger theoretical guarantees than the general results in Dong et al., (2021). We adopt their bounds as a reference baseline, while highlighting the opportunity for future work to establish sharper, model-specific results. This positions our empirical study on firm theoretical footing while opening avenues for more refined analysis.
Remark 2.
The proof techniques used here such as bounding regret via sequential Rademacher complexity and the resulting sample-complexity corollary for ViOlin are direct instantiations of general tools from the bandit literature. What is distinctive in this result is their adaptation to mechanistic nonlinear models (e.g., Mitscherlich, Michaelis-Menten), which integrates domain knowledge with statistical guarantees. This integration ensures that theoretical results are not only mathematically sound but also directly interpretable and actionable in agronomic applications.
@@
Most bandit algorithms rely on the assumption of a well-specified model (or “realizability”), that is, the existence of a function $f^{*}\in\mathcal{F}$ such that the mean reward is $\mu(x)=f^{*}(x)$ for all actions $x$ (Chu et al.,, 2011; Abbasi-Yadkori et al.,, 2011; Agarwal et al.,, 2012; Foster et al.,, 2018). However, in practice, exact realizability rarely holds, and it is important to consider the effect of model misspecification.
Recent work (see, e.g., (Greenewald et al.,, 2021; Lattimore et al.,, 2020; Ghosh et al.,, 2017)) has studied bandits under various notions of misspecification. A common formulation is the uniform $\varepsilon$-misspecification setting, where
@@
For linear bandits, Lattimore et al., (2020) show that the cumulative regret must satisfy
@@
5 Simulation study & Results: Emulating real-world agricultural decision-making
@@ 5.1 From yield to profit: Economic optimization of fertilizer
    Model-based $\epsilon$-greedy algorithm (Algorithm 5): This method selects a random fertilizer rate with probability $\epsilon_{t}=t^{-a}$, with $a$ chosen from $\{0.5,1,1.5\}$, for exploration. With the remaining $(1-\epsilon_{t})$ probability, it chooses the rate closest to the current profit-maximizing estimate $x^{*}$, computed using the fitted model and closed-form expressions where available.
@@
Although these methods differ in their exploration strategies, they all share the same structure: fitting a nonlinear model to guide adaptive decision-making under uncertainty. Accurate initialization of nonlinear model parameters is crucial to the success of these algorithms, especially in small-sample settings. To benchmark the performance of nonlinear model-based strategies, we compare them against two widely used baselines: LinUCB (Chu et al.,, 2011) and $k$-Nearest Neighbor-UCB (kNN-UCB) (Reeve et al.,, 2018).
@@
Table 2: Closed-form expressions for profit-maximizing fertilizer doses under different nonlinear yield models. These highlight how optimal input use depends jointly on agronomic and economic parameters, providing interpretable decision rules.
@@ 5.2 Well-specified setting
In the well-specified setting, data were generated according to a quadratic plateau model of the form:
@@
We evaluated five algorithms: model-based $\epsilon$-greedy, nonlinear UCB, ViOlin (curvature-matched nonlinear bandit), LinUCB (linear model-based UCB), and kNN-UCB (nonparametric UCB using $k=3$ nearest neighbors). Hyperparameters for each of the algorithms are chosen based on theoretical guidelines and grid searches over multiple runs. We choose $\epsilon_{t}=t^{-1.5}$ for $\epsilon$-greedy in this setting, $\alpha=1$ in UCB, linUCB and kNN-UCB for uniformity and fairness in comparison, $\alpha_{1}=2.0,\alpha_{2}=640$ for the ViOlin algorithm based on guidelines in Dong et al., (2021). For each algorithm and parameter configuration, we performed $10$ independent simulation replicates of $T=30$ rounds. Each algorithm fits its specified model to the observed data and selects actions to maximize estimated profit. The regret at each round is defined as the difference between the average profit for the optimal arm (computed using knowledge of the true reward function) and the average profit realized by the arm selected by the algorithm, that is,
@@
where $x^{*}$ denotes the arm maximizing the true profit function, $\theta^{*}$ the true model parameters and $x_{t}$ is the arm selected at round $t$ by the bandit algorithm employed to make decisions. This definition of regret translates directly into wasted profit and fertilizer input, making it a practical metric for resource-efficient agricultural decision-making. For each run, we record cumulative regret, average profit, and select fertilizer rates. In order to quantify the uncertainty, we also plot the distributions of cumulative regret and average profit over 10 replications at round 10, 20, and 30, respectively.
@@
Figure 1 presents the average cumulative regret and the average profit per round for each algorithm. Figure 2 presents boxplots of the cumulative regret and average profit for each algorithm, evaluated at rounds 10, 20, and 30, respectively, across 10 independent replications. Several trends are immediately apparent:
@@
  • Uncertainty Quantification: Figure 2 highlights important differences between model-based approaches (e.g., ViOlin, model based UCB) against linear and nonparametric benchmarks such as lin-UCB and kNN-UCB, particularly in the small-sample regime relevant to agricultural field trials. While kNN-UCB exhibits relatively low variability across replicates, suggesting stable short-term performance, this stability arises primarily from smoothing rather than from capturing the underlying input-response mechanism. As a result, kNN-UCB can suffer from persistent bias, leading to consistently suboptimal mean performance. This is evident in the profit distributions, i.e., the mean profit for kNN-UCB consistently lies below the lower quartile of the nonlinear model-based approaches. In other words, the algorithm produces profits that are reliably clustered, but around a lower and biased center. This suggests that while its variability is small, its systematic bias leads to consistently suboptimal performance. In contrast, nonlinear model-based algorithms exhibit higher run-to-run variability due to parameter estimation uncertainty, particularly in early rounds when data are scarce. However, these algorithms leverage structural assumptions that align with biological processes, enabling faster convergence to the true optimum as more data accrue.
Similar trends were observed across other values of the fertilizer price $p_{x}$ and for all nonlinear response models considered. Taken together, these simulation studies demonstrate a consistent advantage of nonlinear model-based bandit algorithms over both nonparametric and linear parametric alternatives in small-sample regimes. Based on these findings, we subsequently focus our comparisons on the family of nonlinear model-based algorithms, investigating their relative strengths under a range of simulation scenarios to clarify which approaches are most advantageous in which settings.
Figure 1: Average cumulative regret and profit trajectories across 10 replications under the quadratic-plateau model ($p_{x}=0.7$). Nonlinear model-based algorithms consistently outperform linear and nonparametric baselines.
Figure 2: Uncertainty quantification in different Bandit algorithms for profit maximization/regret minimization showing how model-based methods more quickly converge to cost-efficient fertilizer recommendations.
@@
In this next part of the simulation study, we focus on illustrating the implications of using the three non-linear model-based bandit algorithms for learning fertilizer rates sequentially with the goal of profit maximization. Again in a small-sample regime, we run each of our algorithms for $T=30$ rounds and replicate each run 10 times. Since our algorithms choose arms over time, it is important to visualize how the arm choices evolve over time for each of the algorithms. In Figure 3(a), we plot the running proportion of fertilizer rate ranging from $\{0,50,\ldots,250\}$ lb N/ac, selected at each round by each of the algorithm over the 10 replicates for the quadratic plateau model for $p_{x}=\$0.5$/lb N to mimic realistic prices per unit nitrogen fertilizer (Urea) in the Midwest US. Note that in the beginning, the UCB and $\epsilon$-greedy ($\epsilon_{t}=t^{-1.5}$) explore other arms such as $x=100,150$ lb N/ac, but just after a few rounds of exploration, learn the better arm to be $x=200$ lb N/ac. ViOlin is more greedy from the beginning itself as can be seen by the proportions of other arms (other than purple) selected in the initial rounds. In Figure 3(b), in order to assess the affect of increasing fertilizer prices on decision-making for profit maximization, we also plotted the average selected fertilizer rate over time for the three algorithms. We note that the optimum arm (fertilizer rate) choice decreases as $p_{x}$ (price per unit) increases from \$0.3 to \$0.7/lb N. Additionally, since our decisions depend significantly on how well our parameters for the non-linear model are estimated over time, in Figure 4, we also plot the parameter trajectories over time for the quadratic plateau model with $p_{x}=\$0.5$/lb N. Note that all the three algorithms, as data accumulates, the parameter estimates stabilize close to the true parameter values that were used to generate the data. In terms of interpretability, one can, for example, look at the parameter estimates for $x_{0}$ (bottom right), which denotes the threshold value at which the quadratic model plateaus, thus reflecting the saturating point over rounds (or seasons), beyond which additional fertilizer amount provide little benefit. In our simulations, we also examined the sensitivity of parameter estimation to initialization. Because nonlinear least-squares fitting is iterative, the choice of starting values can influence convergence, particularly in small-sample regimes. To reflect realistic practice, we initialized parameters at values close to but not equal to the true generating parameters (see Section D in the Appendix for explicit values used in each model). We observed that biologically plausible initializations led to faster and more stable convergence, while poor starting values could slow parameter recovery. In practice, we recommend using domain knowledge (e.g., agronomic ranges for maximum yield, response rates, or plateau points) to set initial values when fitting nonlinear models. We conducted similar experiments for the other non-linear models and those results are presented in the Appendix.
Figure 4: Parameter trajectories for the quadratic plateau model over rounds